(a^2-a)^2-8(a^2-a)+20
1
Expert's answer
2017-06-05T09:18:09-0400
We use the formula (b-c)^2=b^2-2bc+c^2. In our case we have b=a^2,c=a and we get
(a^2-a)^2-8(a^2-a)+20=a^4-2a^3+a^2-8a^2+8a+20=a^4-2a^3-7a^2+8a+20
Answer a^4-2a^3-7a^2+8a+20